Document Class refman for L^{A}TEX version 2e^{∗}

Copyright (C) 1988 by Hubert Partl Copyright (C) 1994-94 by Axel Kielhorn

1995/03/19

Contents

∗This file has version number v1.9f, last revised 1995/03/19.